a magnetism field of 2X10^3 Am-1 produces a magnetic flux density of 8 phi Tesla in an iron rod. the realative permeability of the rod will be
10^2
10^0
10^1
10^4

Grade:12th pass

1 Answers

Khimraj
5 years ago
relative permeability of rod is given by
ur = B/(uo*H)
H = 2*103 A/m
u0 = 4\pi*10-7
B = 8\pi
so,
ur = 104.
Hope it clears. If you like anser then please approve it.